Two blocks of mass 10 kg and 20 kg moving at speed of 10 m/sec and 20 m/sec respectively in opposite directions approach each other and collide. if the collision is completely inelastic. find the thermal energy in the process

Grade:11

1 Answers

Arun
6 years ago
Dear Student

 

The final velocity of the combination is v (let) towards the initial direction of the 20 kg body Hence the Eq of the conservation of momentum is

(20*20)-(10*10)= 30v

Hence v=10m/z

From the conservation of the total energy we get,

1/2(20*20²+10*10²)=H+1/2(30*10²)

Solving we get H=2000 J

=2 kJ

 

Regards

Arun (askIITians forum expert)